# Rethinking Improved Privacy-Utility Trade-off with Pre-existing Knowledge for DP Training

### v1

@@ IV-B Overview of DP Heterogeneous SGD
Motivated by DP-SGD, we explore an instantiation of $\mathsf{DP\text{-}Hero}$ to generate heterogeneous noise and then add a “wisdom” (guided by prior learned knowledge) heterogeneous noise. Accordingly, we instantiate DP-SGD [14] as the basis and replace its i.i.d. noise with heterogeneous noise. In DP-SGD, the standard deviation $\sigma$ of $\mathcal{N}(0,\sigma^{2})$ is constant for each layer; however, our mechanism guided by $\mathsf{SVec}$ adds different noise vectors for model updates at each iteration. With $\mathsf{SVec}^{(t-1)}$, the added noise to each layer is guided by the learned model in the aspects of scales and noise space at every iteration.
Using $\mathsf{DP\text{-}Hero}$ SGD, we implement an instantiated scheme of training a model starting from random initialization. The first step is generating heterogeneous noise building on the covariance matrix of the model. By principal component analysis (PCA) [49], the noise matrix is tuned via the covariance matrix, which aligns with the subspace in which features exist. When training with SGD, updatable gradients computed in the backpropagation are added by noise, whose scales are guided by the generated subspace. We consider extracting pre-existing knowledge from whole model parameters rather than a layer to capture the whole statistical space. In this way, the noise space is more comprehensive, and the noise scale is more adaptive to the feature space.
IV-C Detailed Construction
IV-C1 Construction of $\mathsf{DP\text{-}Hero}$ SGD
For clarification, we explain the step-by-step construction of $\mathsf{DP\text{-}Hero}$ SGD.
@@ IV-C1 Construction of $\mathsf{DP\text{-}Hero}$ SGD
Step-2. In our implementation, $\mathsf{SVec}^{(t-1)}$ can be realized by following Algorithm 2 using $\mathbf{W}^{(t-1)}$. Since $\mathsf{SVec}^{(t-1)}$ is varied at each training iteration, $\mathsf{SVec}^{(t-1)}$-guided noise distribution operating on gradients is varied during the whole training process. $\mathsf{SVec}^{(t-1)}$ contains the computed sub-space $\mathbf{B}^{(t-1)}$ and eigenvalues matrix $\mathbf{V}^{(t-1)}$ extracted from prior-learned model. From a practical view, $\mathbf{B}^{(t-1)}$ configures the direction of the noise to be added. $\mathbf{V}^{(t-1)}$ generated from singular value decomposition is utilized to scale the noise distribution. Here, independent and identically distributed noise can be sampled from a standard noise distribution $\mathcal{N}$, such as Gaussian and Laplace distributions. The generation of $\mathsf{SVec}^{(t-1)}$ does not introduce extra leakage since $\mathbf{W}^{(t-1)}$ learned in the prior $t-1$ iterations has been well-protected through $\mathsf{DP\text{-}Hero}$ SGD.
@@
IV-C2 Construction of Noise Guidance
The math tool, principal component analysis (PCA) [50] performs analyzing data represented by inter-correlated quantitative dependent variables. It forms a set of new orthogonal variables, called components, depending on the matrix eigen-decomposition and singular value decomposition (SVD). Given a matrix $\mathbf{X}$, of column-wise mean equal to 00, the multiplication $\mathbf{X}^{\top}\mathbf{X}$ is a correlation matrix. Later, a diagonal matrix of the (non-zero) eigenvalues of $\mathbf{X}^{\top}\mathbf{X}$ is extracted together with the eigenvectors. Essentially, PCA simplifies data representation and decomposes its corresponding structures.
We propose a simple yet efficient approach by examining the model parameters as a result of knowledge integration over diverse features extracted from private data. As in Algorithm 2, we employ the PCA decomposition [49] to extract knowledge learned by the training model and apply generated guidance $\mathsf{SVec}^{(t)}$ at iteration $t$ to adjust noise addition at the next iteration. PCA decomposition can extract knowledge from representative data (i.e., model parameters in our setting) by analyzing inter-correlated quantitative dependence. Normally, a neural network kernel extracting the features from the images is a matrix that moves over the input data to perform the dot product with the sub-region of input data. Denote $\mathbb{R}$ to be the real number set. Let $\mathbf{b}=[b_{1},b_{2},\ldots,b_{k}]$ be a vector, and $\mathbf{B}=[\mathbf{b_{1}},\mathbf{b_{2}},\ldots,\mathbf{b_{d}}]^{\top}\in\mathbb{R}^{n\times m}$ be a matrix.
@@ IV-C2 Construction of Noise Guidance
IV-C3 Noise Guidance through Pre-existing Knowledge
@@
For a non-private model, $\mathbf{W}$ converges to a stable status through uncountable routes of optimizing model parameters. Noise addition becomes complicated if referring to different optimizing tools; it is not generic anymore. DP-SGD sets a fixed noise scale at different training iterations. Noise addition on $\mathbf{W}$ inevitably has a negative contribution to extracting features over private data compared with pure parameters. By rethinking DP training from sketch (i.e., random to convergence), varying
@@
$\mathsf{SVec}$ achieves improved allocation of parameter-wise heterogeneous noise at each training iteration with the constraint of a preset privacy budget. Such an automatic allocation is generated from the prior-iteration evaluation of the training model in a differentially private manner. From this viewpoint, injecting noise into the model parameters contributes negatively to both the knowledge and the process of knowledge integration. Compared with DP-SGD, the proposed method mitigates destroying the process of knowledge integration while keeping the learned knowledge unchanged. Different grid search of tuning hyper-parameters, $\mathsf{DP\text{-}Hero}$ SGD adjusts the intermediate training process via instantly learnable parameters rather than setting a set of possibilities. Combining grid search (vertically tuning) and $\mathsf{DP\text{-}Hero}$ SGD (horizontally tuning) may further boost the automatic optimization of DP learning in an algorithmic view.
IV-D Privacy Analysis and Theoretical Explanation
We first analyze the DP guarantee of $\mathsf{DP\text{-}Hero}$ SGD, which provides identical protection as standard DP-SGD as shown in Theorem 2. Then, building on Theorem 2, we instantiate $\mathsf{DP\text{-}Hero}$ SGD regarding the parameters configuration as Theorem 3.
@@ IV-D Privacy Analysis and Theoretical Explanation
Let a random mechanism $\mathcal{M}^{(t)}$ be $(\epsilon^{(t)},\delta)$-differential privacy at the iteration $t$. A $\mathsf{DP\text{-}Hero}$ mechanism ~$\tilde{\mathcal{M}}^{(t)}$ parameterized by $(\tilde{\epsilon}^{(t)},\delta)$ is $({\epsilon},\delta)$-differential privacy if $\tilde{\sigma}=\sigma$.
@@
Standard DP-SGD is $(\epsilon,\delta)$-differentially private if $\sigma\geq c_{2}q\sqrt{T\log(1/\delta)}/{\epsilon}$ for any $\delta>0$ [14]. The $q,T$ are, respectively, sampling probability and the number of steps relevant to model training. The $c$ is a constant for all DP mechanisms. Take $\mathcal{M}^{({t})}$ to be a $\mathsf{DP\text{-}Hero}$ random mechanism that is derived from $(\epsilon,\delta)$-differential privacy. The $\mathcal{M}_{\mathsf{t}}$ has the same configuration of $q,T,c$ due to the identical training procedure. If $\sigma$ is unchanged, $\mathcal{M}^{({t})}$ also satisfies $\sigma\geq c_{2}{q\sqrt{T\log(1/\delta)}}/{\epsilon}$ for any $\delta>0$. Thus, $\mathcal{M}_{\mathsf{t}}$ is $(\epsilon,\delta)$-differentially private. ∎
@@
$\mathsf{DP\text{-}Hero}$ SGD parameterized by $\mathcal{N}(0,\tilde{\sigma}^{2})$ and standard DP-SGD parameterized by $\mathcal{N}(0,\sigma^{2})$ satisfy $\tilde{\sigma}=\sigma$ such that the $i$-th entry of diagonal matrix $\mathsf{SVec}$ is set to be, $v_{i}={\mathsf{egv}_{i}\cdot\sqrt{k}\cdot\sigma}/{\sqrt{\sum_{i=1}^{k}{\mathsf{egv}_{i}}^{2}}}$.
@@
Building on $\alpha$-Rényi divergence and privacy loss, concentrated differential privacy (CDP) [27] allows improved computation mitigating single-query loss and high probability bounds for accurately analyzing the cumulative loss. It centralizes privacy loss around zero, maintaining sub-Gaussian characteristics that make larger deviations from zero increasingly improbable. In return, zero-CDP implies $(\epsilon_{\rho,\delta},\delta)$-DP as restated in Theorem 4 [27].
Definition 8 (zero-CDP [27]).
@@
Theorem 4 (From zero-CDP to $(\epsilon,\delta)$-DP [27]).
@@
At last, since we have aligned the privacy guarantee of $\mathsf{DP\text{-}Hero}$ with the standard DP, we follow the standard composition-paradigm proof [28] under the definition of zCDP [51, 27, 16] through Rényi Divergence by Bun et al. [27] for a tight analysis, as shown in Theorem 5.
@@
[…]
@@
By Definition 3 and $\mathbb{E}\left[e^{(\alpha-1)\mathcal{L}_{\mathsf{Pri}}^{(o),(t)}}\right]$ [16], we have,
@@
Subject to $\sigma={\sqrt{2T\log{(1/\delta)}}}/{\epsilon}$, we use $\alpha=\frac{\epsilon+(\max_{t}\rho^{(t)})}{2\cdot(\max_{t}\rho^{(t)})}$ as derived in [16], and compute,
@@
still satisfies original DP definition, as in [27, 28]. ∎
IV-E Linear Layer Analysis as an Example
@@
[…]
Consider that noise variable $n$ being injected into each feature could be continuous ideally. Since it is sampled from a distribution with a mean value of 00, the integration of $n$ equals 00, which could be removed for simplification.
@@
$(\mathbf{W}+\mathbf{N})^{\top}x_{i}{y_{i}}^{\prime}-\mathbf{W}^{\top}x_{i}y_{i}=(\mathbf{W}+\mathbf{N})^{\top}x_{i}{(y_{i}+\Delta y_{i})}-\mathbf{W}^{\top}x_{i}y_{i}$ (19)
$=\mathbf{W}^{\top}x_{i}{\Delta y_{i}}+\mathbf{N}^{\top}x_{i}{y_{i}}+\mathbf{N}^{\top}x_{i}{\Delta y_{i}}=(\mathbf{W}+\mathbf{N})^{\top}x_{i}{\Delta y_{i}}+\mathbf{N}^{\top}x_{i}{y_{i}}$
@@
Specifically, if $(\mathbf{W}+\mathbf{N})^{\top}x_{i}$ is close to $y_{i}$, the differentially-private (noisy for short) model accuracy is high. To attain the minimizer, we could solve Equation 19 by $\mathbf{W}\bot\mathbf{N}$. In this example analysis, attaining support for the noise-model relation is enough for the initial exploration.
@@
V Experimental Evaluation and Explanation
Our experiments are conducted on a commodity PC running Ubuntu with Intel Xeon(R) E5-2630 v3 CPU, 31.3 GiB RAM, and GeForce RTX $1080$ Ti GPU. In this section, we report the convergence/training performance and test accuracy (varying with $\epsilon$) by conducting an extensive comparison with state-of-the-arts [14, 52, 53, 54, 55, 16, 56, 46] over standard benchmark datasets. By employing GridCam [48], we visualize differentially private training to show the difference in representation.
V-A Experimental Setup
V-A1 Configuration and Dataset
The baseline DP-SGD implementation is pyvacy (https://github.com/ChrisWaites/pyvacy). We configure experimental parameters with reference to [14]’s setting. To be specific, we configure lot size $L=10,50,200,400$, $\delta=1.0^{-5}\text{ or }1.0^{-6}$, and learning rate $\eta=0.1\text{ or }0.2$. The noise level $\sigma$ is set to be $0.5,1,3,5,7,10$ for comprehensive comparison. Fairly, we use identical $\epsilon$ as in state-of-the-art and compare test accuracy.
Experimental evaluations are performed on the MNIST dataset [57] and the CIFAR-10 dataset [58]. MNIST dataset includes $10$ classes of hand-written digits of $28\times 28$ gray-scale. It contains $60,000$ training examples and $10,000$ testing examples. CIFAR-10 dataset contains $10$ classes of images, of $32\times 32$ color-scale with three channels, It contains $50,000$ in training examples and $10,000$ in testing examples.
V-A2 Model Architecture
On the MNIST dataset, we use LeNet [57], which reaches accuracy of $99\%$ in about $10$ epochs without privacy. On CIFAR-10, we use two convolutional layers followed by two fully connected layers. In detail, convolution layers use $5\times 5$ convolutions, followed by a ReLU and $2\times 2$ max-pooling. The latter is flattened to a vector that gets fed into two fully connected layers with $384$ units. This architecture, non-privately, can get to about $86\%$ accuracy in $\sim 200$ epochs.
V-B Model Utility and Training Performance
V-B1 Convergence Analysis

### v3

@@ Introduction
Motivated by DP-SGD, we explore an instantiation of $\mathsf{DP\text{-}Hero}$ to generate heterogeneous noise and then add a “wisdom” (guided by prior learned knowledge) heterogeneous noise. Accordingly, we instantiate DP-SGD (Abadi et al. 2016) as the basis and replace its i.i.d. noise with heterogeneous noise. In DP-SGD, the standard deviation $\sigma$ of $\mathcal{N}(0,\sigma^{2})$ is constant for each layer; however, our mechanism guided by $\mathsf{SVec}$ adds different noise vectors for model updates at each iteration. With $\mathsf{SVec}^{(t-1)}$, the added noise to each layer is guided by the learned model in the aspects of scales and noise space at every iteration.
Using $\mathsf{DP\text{-}Hero}$ SGD, we implement an instantiated scheme of training a model starting from random initialization. The first step is generating heterogeneous noise building on the covariance matrix of the model. By principal component analysis (PCA) (Jolliffe and Cadima 2016), the noise matrix is tuned via the covariance matrix, which aligns with the subspace in which features exist. PCA provides a natural mechanism to uncover and exploit heterogeneity in the underlying data distribution. Specifically, PCA rotates the model parameters into a new orthogonal basis where each principal component corresponds to an axis of maximal variance, and the associated eigenvalues quantify the variance along each direction. This decomposition reveals strong anisotropy: some directions exhibit substantially higher variability than others. In the context of DP learning, such heterogeneity is critical, as the sensitivity of the data to perturbations varies across directions. Allocating homogeneous noise across all dimensions fails to respect this structure, disproportionately affecting components with low intrinsic variance. By leveraging PCA, noise can be modulated according to the variance structure, i.e., injecting smaller noise in high-variance directions that are more robust to perturbations, and simultaneously allocating more regularization or larger noise to fragile, low-variance directions.
Formally, by transforming the model parameters into the PCA basis, applying direction-dependent noise proportional to the inverse eigenvalues of the covariance matrix, and transforming back, one achieves a heterogeneous noise model that preserves critical information while satisfying DP constraints. Thus, PCA serves as a powerful tool for constructing heterogeneity-aware noise mechanisms, enabling significantly improved privacy-utility trade-offs compared to uniform noise baselines. When training with SGD, updatable gradients computed in the backpropagation are added by noise, whose scales are guided by the subspace generated by PCA. We consider extracting pre-existing knowledge from whole model parameters rather than a layer to capture the whole statistical space. In this way, the noise space is more comprehensive, and the noise scale is more adaptive to the feature space.
Detailed Construction
Construction of $\mathsf{DP\text{-}Hero}$ SGD
Regarding Definition 7, achieving strong utility under tight privacy constraints demands carefully shaped noise: uniform noise across dimensions can disproportionately damage sensitive components, leading to unnecessary degradation in optimization performance. This motivates the use of PCA, which reveals intrinsic heterogeneity in the update space by decomposing it into orthogonal directions with varying variance. PCA identifies principal directions where the model is naturally more robust to perturbations, enabling noise allocation that is inversely aligned with direction-specific sensitivity. By injecting smaller noise along high-variance directions and larger noise along low-variance ones, a PCA-based noise mechanism better preserves informative structures in the model updates while satisfying the same overall DP guarantees. Thus, PCA provides a principled and structure-aware strategy to optimize the critical trade-off between privacy preservation and utility maximization articulated in Definition 7.
@@
Step-2. In our implementation, $\mathsf{SVec}^{(t-1)}$ can be realized by following Algorithm 2 using $\mathbf{W}^{(t-1)}$. Since $\mathsf{SVec}^{(t-1)}$ is varied at each training iteration, $\mathsf{SVec}^{(t-1)}$-guided noise distribution operating on gradients is varied during the whole training process. $\mathsf{SVec}^{(t-1)}$ contains the computed sub-space $\mathbf{B}^{(t-1)}$ and eigenvalues matrix $\mathbf{V}^{(t-1)}$ extracted from prior-learned model. From a practical view, $\mathbf{B}^{(t-1)}$ configures the direction of the noise to be added. $\mathbf{V}^{(t-1)}$ generated from singular value decomposition is utilized to scale the noise distribution. Here, independent and identically distributed noise can be sampled from a standard noise distribution $\mathcal{N}$, such as Gaussian and Laplace distributions. The generation of $\mathsf{SVec}^{(t-1)}$ does not introduce extra leakage since $\mathbf{W}^{(t-1)}$ learned in the prior t1t-1 iterations has been well-protected through $\mathsf{DP\text{-}Hero}$ SGD.
@@
Construction of Noise Guidance
The math tool, principal component analysis (PCA) (Shawe-Taylor and Williams 2002) performs analyzing data represented by inter-correlated quantitative dependent variables. It forms a set of new orthogonal variables, called components, depending on the matrix eigen-decomposition and singular value decomposition (SVD). Given a matrix $\mathbf{X}$, of column-wise mean equal to 0, the multiplication $\mathbf{X}^{\top}\mathbf{X}$ is a correlation matrix. Later, a diagonal matrix of the (non-zero) eigenvalues of $\mathbf{X}^{\top}\mathbf{X}$ is extracted together with the eigenvectors. Essentially, PCA simplifies data representation and decomposes its corresponding structures.
We propose a simple yet efficient approach by examining the model parameters as a result of knowledge integration over diverse features extracted from private data. As in Algorithm 2, we employ the PCA decomposition (Jolliffe and Cadima 2016) to extract knowledge learned by the training model and apply generated guidance $\mathsf{SVec}^{(t)}$ at iteration $t$ to adjust noise addition at the next iteration. PCA decomposition can extract knowledge from representative data (i.e., model parameters in our setting) by analyzing inter-correlated quantitative dependence. Normally, a neural network kernel extracting the features from the images is a matrix that moves over the input data to perform the dot product with the sub-region of input data. Denote $\mathbb{R}$ to be the real number set. Let $\mathbf{b}=[b_{1},b_{2},\ldots,b_{k}]$ be a vector, and $\mathbf{B}=[\mathbf{b_{1}},\mathbf{b_{2}},\ldots,\mathbf{b_{d}}]^{\top}\in\mathbb{R}^{n\times m}$ be a matrix.
@@
Noise Guidance through Pre-existing Knowledge
@@
For a non-private model, $\mathbf{W}$ converges to a stable status through uncountable routes of optimizing model parameters. Noise addition becomes complicated if we refer to different optimization tools; it is no longer generic. The addition of noise in $\mathbf{W}$ inevitably has a negative contribution to the extraction of features from private data compared to pure parameters.
@@
$\mathsf{SVec}$ achieves improved allocation of parameter-wise heterogeneous noise at each training iteration with the constraint of a preset privacy budget. This automatic allocation is generated from the prioritization evaluation of the training model in a differentially private manner. From this viewpoint, injecting noise into the model parameters negatively contributes to both the knowledge and the process of knowledge integration. Compared with DP-SGD, the proposed method mitigates the destruction of the process of knowledge integration while keeping the learned knowledge unchanged. Different grid search for tuning hyperparameters, $\mathsf{DP\text{-}Hero}$ SGD adjusts the intermediate training process via instantaneous learnable parameters rather than setting a set of possibilities. Combining grid search (vertically tuned) and $\mathsf{DP\text{-}Hero}$ SGD (horizontally tuning) may further boost the automatic optimization of DP learning in an algorithmic view.
Federated Training with $\mathsf{DP\text{-}Hero}$
FedFed (Yang et al. 2023) is a hierarchical federated optimization protocol that divides clients into groups and employs a two-level communication structure. Clients locally compute model updates using stochastic gradient descent (SGD), and then communicate either intra-group or inter-group for aggregations, significantly reducing the global communication cost. To extend FedFed with formal DP guarantees, we apply per-client $\mathsf{DP\text{-}Hero}$ SGD in the local training phase. The modified procedure for each client in FedFed is as Algorithm 3.
The training process proceeds in rounds orchestrated by a central server. At the start, the server distributes the initial global model $\phi^{0}$ and a globally shared dataset $\mathcal{D}^{s}$ to all clients, who each combine it with their local private dataset $\mathcal{D}^{k}$ to form $\mathcal{D}_{t}^{k}=\mathcal{D}^{k}\cup\mathcal{D}^{s}$. In each communication round $r$, the server randomly selects a subset of clients $\mathcal{C}_{r}\subseteq{1,\ldots,K}$, transmits the current global model $\phi^{r}$, and waits while each selected client $k\in\mathcal{C}_{r}$ performs $E_{r}$ epochs of local training using the DP-Hero SGD optimizer on their combined dataset $\mathcal{D}_{t}^{k}$. This approach ensures local updates achieve differential privacy via gradient clipping and noise addition. Upon completion, clients send their updated model parameters $\phi_{k}^{r+1}$ to the server, which aggregates these updates (e.g., via $AGG({\phi_{k}^{r+1}})$) to form the next global model $\phi^{r+1}$. This iterative process allows the system to collaboratively train a privateglobal model with DP guarantee, while benefiting from the scalability and communication efficiency of the FedFed.
Algorithm 3 FedFed with $\mathsf{DP\text{-}Hero}$ SGD
Server Input: initial global model $\phi^{0}$, communication round $T_{r}$.
Client $k$’s Input: local epochs $E_{r}$, local private datasets $\mathcal{D}^{k}$, learning rate $\eta_{k}$.
Initialization: server distributes the initial model $\phi^{0}$ to all clients,
 Generate globally shared dataset $\mathcal{D}^{s}$.
 Distribute $\mathcal{D}^{s}$ to all clients and $\mathcal{D}_{t}^{k}=\mathcal{D}^{k}\cup\mathcal{D}^{s}$.
Server Executes:
for each round $r=1,2,\cdots,T_{r}$ do
  server samples a subset of clients $C_{r}\subseteq\left\{1,...,K\right\}$
  server communicates $\phi^{r}$ to selected clients $k\in C_{r}$
  for  each client $k\in C_{r}$ in parallel  do
   $\phi_{k}^{r+1}\leftarrow\textbf{Client\_Training}(k,\phi^{r}$)
  end for
  $\phi^{r+1}\leftarrow AGG(\phi_{k}^{r+1})$
end for
Client_Training($k,\phi^{r}$):
$\phi^{r}$ initialize local model $\phi_{k}^{r}$
for each local epoch $e$ with $e=1,2,\cdots,E_{r}$ do
  $\phi_{k}^{r+1}\leftarrow$ $\mathsf{DP\text{-}Hero}$ SGD update with $\mathcal{D}_{t}^{k}$
end for
Return $\phi_{k}^{r+1}$ to server
Privacy Analysis and Theoretical Explanation
We establish the fundamental privacy guarantee for DP-Hero mechanisms. It shows that if the noise scale ($\sigma$) is chosen according to the lower bound established by  (Abadi et al. 2016), then the $\mathsf{DP\text{-}Hero}$ mechanism at each iteration attains ($\epsilon,\delta$)-DP. Theorem 2 ensures that the per-iteration privacy risk of DP-Hero is comparable to that of classical DP mechanisms, provided the noise is correctly calibrated.
@@
Let a random mechanism $\mathcal{M}^{(t)}$ be $(\epsilon^{(t)},\delta)$-differential privacy at the iteration $t$. A $\mathsf{DP\text{-}Hero}$ mechanism $\tilde{\mathcal{M}}^{(t)}$ parameterized by $(\tilde{\epsilon}^{(t)},\delta)$ is $({\epsilon},\delta)$-differential privacy if $\tilde{\sigma}=\sigma$, where $\sigma\geq c_{2}q\sqrt{T\log(1/\delta)}/{\epsilon}$ (Abadi et al. 2016).
@@
Standard DP-SGD is $(\epsilon,\delta)$-differentially private if $\sigma\geq c_{2}q\sqrt{T\log(1/\delta)}/{\epsilon}$ for any $\delta>0$ (Abadi et al. 2016). The $q,T$ are, respectively, sampling probability and the number of steps relevant to model training. The $c$ is a constant for all DP mechanisms. Take $\mathcal{M}^{({t})}$ to be a $\mathsf{DP\text{-}Hero}$ random mechanism that is derived from $(\epsilon,\delta)$-differential privacy. The $\mathcal{M}_{\mathsf{t}}$ has the same configuration of $q,T,c$ due to the identical training procedure. If $\sigma$ is unchanged, $\mathcal{M}^{({t})}$ also satisfies $\sigma\geq c_{2}{q\sqrt{T\log(1/\delta)}}/{\epsilon}$ for any $\delta>0$. Thus, $\mathcal{M}_{\mathsf{t}}$ is $(\epsilon,\delta)$-differentially private. ∎
Then, we demonstrate that by appropriately parameterizing the diagonal scaling matrix ($\mathsf{SVec}$), the total variance of noise injected by $\mathsf{DP\text{-}Hero}$ SGD can be made equal to that of standard DP-SGD. This means that the two mechanisms have matching privacy and utility properties from the perspective of total noise magnitude as in Theorem 3.
@@
Let $\mathsf{DP\text{-}Hero}$ SGD be parameterized by $\mathcal{N}(0,\tilde{\sigma}^{2})$ and standard DP-SGD be parameterized by $\mathcal{N}(0,\sigma^{2})$, repectively. Consider $\mathsf{DP\text{-}Hero}$ SGD adds noise from $\mathcal{N}(0,\tilde{\sigma}^{2})$ and standard DP-SGD uses $\mathcal{N}(0,\sigma^{2})$. Let $\mathsf{SVec}$ be a diagonal matrix whose $i$-th diagonal entry is defined as $v_{i}={\mathsf{egv}_{i}\cdot\sqrt{k}\cdot\sigma}/{\sqrt{\sum_{i=1}^{k}{\mathsf{egv}_{i}}^{2}}}$, where $k$ denotes the dimension. Then, if $\tilde{\sigma}=\sigma$, the total variance of the noise injected by DP-Hero SGD with $\mathsf{SVec}$ equals that of standard DP-SGD.
@@
Building on $\alpha$-Rényi divergence and privacy loss, concentrated differential privacy (CDP) (Bun and Steinke 2016) allows improved computation mitigating single-query loss and high probability bounds for accurately analyzing the cumulative loss. It centralizes privacy loss around zero, maintaining sub-Gaussian characteristics that make larger deviations from zero increasingly improbable. In return, zero-CDP implies $(\epsilon_{\rho,\delta},\delta)$-DP as restated in Theorem 4 (Bun and Steinke 2016).
Definition 8 (zero-CDP (Bun and Steinke 2016)).
@@
Theorem 4 (From zero-CDP to $(\epsilon,\delta)$-DP (Bun and Steinke 2016)).
@@
At last, since we have aligned the privacy guarantee of $\mathsf{DP\text{-}Hero}$ with the standard DP, we follow the standard composition-paradigm proof (Mironov 2017) under the definition of zCDP (Dwork and Rothblum 2016; Bun and Steinke 2016; Yu et al. 2019) through Rényi Divergence by Bun et al. (Bun and Steinke 2016) for a tight analysis, resulting in Theorem 5.
@@
$=\int_{\mathbb{R}}e^{\left({-\alpha(x-\mathcal{F}(D))^{2}}/{2\sigma^{2}}-{(1-\alpha)(x-\mathcal{F}(D^{\prime}))^{2}}/{2\sigma^{2}}\right)}dx$
$=\int_{\mathbb{R}}e^{\left(-{(x-(\alpha\mathcal{F}(D)+(1-\alpha)\mathcal{F}(D^{\prime})))^{2}}/{2\sigma^{2}}\right)}dx$
$\quad+\int_{\mathbb{R}}e^{\left({(\alpha\mathcal{F}(D)+(1-\alpha)\mathcal{F}(D^{\prime}))^{2}-\alpha\mathcal{F}(D)^{2}}/{2\sigma^{2}}\right)}dx$
$\quad-\int_{\mathbb{R}}e^{\left({(1-\alpha)\mathcal{F}(D^{\prime})^{2}}/{2\sigma^{2}}\right)}dx$
@@
By Definition 3 and $\mathbb{E}\left[e^{(\alpha-1)\mathcal{L}_{\mathsf{Pri}}^{(o),(t)}}\right]$ (Yu et al. 2019), we have,
@@
Subject to $\sigma={\sqrt{2T\log{(1/\delta)}}}/{\epsilon}$, we use $\alpha=\frac{\epsilon+(\max_{t}\rho^{(t)})}{2\cdot(\max_{t}\rho^{(t)})}$ as derived in (Yu et al. 2019), and compute,
@@
still satisfies original DP definition, as in (Bun and Steinke 2016; Mironov 2017). ∎
Together, these theorems ensure that DP-Hero SGD matches standard practice in DP-SGD both in per-step and composed privacy, and that the parameterization of $\mathsf{DP\text{-}Hero}$ SGD can be mapped to the well-established privacy analysis in the previous work.
Linear Layer Analysis as an Example
@@
$\arg\min_{\|\mathbf{N}\|}|\sum_{i}\int(\mathbf{W}+\mathbf{N})x_{i}{y_{i}}^{\prime}\,dn-\sum_{i}\int\mathbf{W}x_{i}y_{i}\,dn|$ (17)
Consider that noise variable $n$ being injected into each feature could be continuous ideally. Since it is sampled from a distribution with a mean value of 0, the integration of $n$ equals 0, which could be removed for simplification.
@@
$(\mathbf{W}+\mathbf{N})^{\top}x_{i}{y_{i}}^{\prime}-\mathbf{W}^{\top}x_{i}y_{i}$ (20)
$=(\mathbf{W}+\mathbf{N})^{\top}x_{i}{(y_{i}+\Delta y_{i})}-\mathbf{W}^{\top}x_{i}y_{i}$
$=\mathbf{W}^{\top}x_{i}{\Delta y_{i}}+\mathbf{N}^{\top}x_{i}{y_{i}}+\mathbf{N}^{\top}x_{i}{\Delta y_{i}}$
$=(\mathbf{W}+\mathbf{N})^{\top}x_{i}{\Delta y_{i}}+\mathbf{N}^{\top}x_{i}{y_{i}}$
@@
Specifically, if $(\mathbf{W}+\mathbf{N})^{\top}x_{i}$ is close to $y_{i}$, the differentially-private (noisy for short) model accuracy is high. To attain the minimizer, we could solve Equation 20 by $\mathbf{W}\bot\mathbf{N}$. In this example analysis, attaining support for the noise-model relation is enough for simplification.
@@
Experimental Evaluation and Explanation
Our experiments are conducted on a commodity PC running Ubuntu with Intel Xeon(R) E5-2630 v3 CPU, 31.3 GiB RAM, and GeForce RTX $3090$ Ti GPU. In this section, we report the convergence/training performance and test accuracy (varying with $\epsilon$) by conducting an extensive comparison with state-of-the-arts, including DP-SGD (Abadi et al. 2016), IPA-RF (Feldman and Zrnic 2021), GDP (Bu et al. 2019), SAS-DP (Chen and Lee 2020), DP-GED (Nasr, Shokri, and Houmansadr 2020), DP-MP (Yu et al. 2019), LDP (Ghazi et al. 2021), FeatureDP (Tramèr and Boneh 2021), and FedFed (Yang et al. 2023) over standard benchmark datasets. By employing GridCam (Selvaraju et al. 2017), we visualize differentially private training to show the difference in representation.
Experimental Setup
Configuration and Dataset
The baseline DP-SGD implementation is pyvacy (https://github.com/ChrisWaites/pyvacy), and , while federated learning with $\mathsf{DP\text{-}Hero}$ has been realized in FedFed framework (Yang et al. 2023). We configure experimental parameters with reference to (Abadi et al. 2016)’s setting. To be specific, we configure lot size $L=10,50,200,400$, $\delta=1.0^{-5}\text{ or }1.0^{-6}$, and learning rate $\eta=0.1\text{ or }0.2$. The noise level $\sigma$ is set to be $0.5,1,3,5,7,10$ for comprehensive comparison. Fairly, we use identical $\epsilon$ as in state-of-the-art and compare test accuracy.
Experimental evaluations are performed on the MNIST dataset (LeCun et al. 1998) and the CIFAR-10 dataset (Krizhevsky, Hinton et al. 2009). MNIST dataset includes $10$ classes of hand-written digits of $28\times 28$ gray-scale. It contains $60,000$ training examples and $10,000$ testing examples. CIFAR-10 dataset contains $10$ classes of images, of $32\times 32$ color-scale with three channels, It contains $50,000$ in training examples and $10,000$ in testing examples.
Model Architecture
On the MNIST dataset, we use LeNet (LeCun et al. 1998), which reaches accuracy of $99\%$ in about $10$ epochs without privacy. On CIFAR-10, we use two convolutional layers followed by two fully connected layers. In detail, convolution layers use $5\times 5$ convolutions, followed by a ReLU and $2\times 2$ max-pooling. The latter is flattened to a vector that gets fed into two fully connected layers with $384$ units. This architecture, non-privately, can get to about $86\%$ accuracy in $\sim 200$ epochs.
Model Utility and Training Performance
Convergence Analysis